OpenStudy (anonymous):

****** metal ************* help please the altitude to the hypotenuse of a right triangle divides the hypotenuse into segment that are 2cm long and 5cm long. find the length of the altitude to the nearest tenth of a percent

OpenStudy (anonymous):

diagram

jimthompson5910 (jim_thompson5910):

this is a word problem that's just a translation of what you've already done

jimthompson5910 (jim_thompson5910):

2/x = x/5 2*5 = x*x keep going to solve for x

OpenStudy (anonymous):

need the answer compleate

jimthompson5910 (jim_thompson5910):

try it out

OpenStudy (anonymous):

sorry i cant do it i suck at this

jimthompson5910 (jim_thompson5910):

give it a shot

jimthompson5910 (jim_thompson5910):

its ok if you don't get it, that's what practice is for

OpenStudy (anonymous):

i cnt sorry please just help me

jimthompson5910 (jim_thompson5910):

2*5 = ____ fill in the blank

OpenStudy (anonymous):

x\

jimthompson5910 (jim_thompson5910):

2 times 5 = _____

jimthompson5910 (jim_thompson5910):

2*5 means 2 times 5

OpenStudy (anonymous):

10

jimthompson5910 (jim_thompson5910):

x times x = ______

OpenStudy (anonymous):

2x

jimthompson5910 (jim_thompson5910):

that's x+x

jimthompson5910 (jim_thompson5910):

x times x is x^2

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

this means 2*5 = x*x turns into 10 = x^2 x^2 = 10 so x = _____

jimthompson5910 (jim_thompson5910):

it's not 5

OpenStudy (anonymous):

100

jimthompson5910 (jim_thompson5910):

it's not 100

jimthompson5910 (jim_thompson5910):

undo the square by taking the square root of both sides

OpenStudy (anonymous):

3.16

jimthompson5910 (jim_thompson5910):

round that to the nearest tenth

OpenStudy (anonymous):

3 or 3.1

jimthompson5910 (jim_thompson5910):

what number is in the tenths place

OpenStudy (anonymous):

idk 3

jimthompson5910 (jim_thompson5910):

no, that's in the units place

OpenStudy (anonymous):

mmm i give up i have noo idea

jimthompson5910 (jim_thompson5910):

hint: hundreds tens units . tenths hundredths

OpenStudy (anonymous):

??? im lost sorry :'(

jimthompson5910 (jim_thompson5910):

here's an example http://cstl.syr.edu/fipse/decunit/intrdec/introgif/Fig1.gif

jimthompson5910 (jim_thompson5910):

they are using the number 127.5369

OpenStudy (anonymous):

oki got it can you help me #43

OpenStudy (anonymous):

is it 3.16227

jimthompson5910 (jim_thompson5910):

no it's 3.2

OpenStudy (anonymous):

wow my friend dum

OpenStudy (anonymous):

can you help me on #43 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

check out this link http://www.math.nmsu.edu/~breakingaway/Lessons/PTUST/PTUST.html